### v1

# Target Aggregate Data Adjustment Method for Transportability Analysis Utilizing Summary-Level Data from the Target Population

Abstract
Transportability analysis is a causal inference framework used to evaluate the external validity of randomized clinical trials (RCTs) or observational studies. Most existing transportability analysis methods require individual patient-level data (IPD) for both the source and the target population, narrowing its applicability when only target aggregate-level data (AgD) is available. Besides, accounting for censoring is essential to reduce bias in longitudinal data, yet AgD-based transportability methods in the presence of censoring remain underexplored. Here, we propose a two-stage weighting framework named “Target Aggregate Data Adjustment” (TADA) to address the mentioned challenges simultaneously. TADA is designed as a two-stage weighting scheme to simultaneously adjust for both censoring bias and distributional imbalances of effect modifiers (EM), where the final weights are the product of the inverse probability of censoring weights and participation weights derived using the method of moments. We have conducted an extensive simulation study to evaluate TADA’s performance. Our results indicate that TADA can effectively control the bias resulting from censoring within a non-extreme range suitable for most practical scenarios, and enhance the application and clinical interpretability of transportability analyses in settings with limited data availability.
Keywords transportability analysis  $\cdot$ causal inference  $\cdot$ survival analysis  $\cdot$ aggregate-level data  $\cdot$ inverse probability of censoring weights  $\cdot$ method of moments
1 Introduction
Transportability analysis is a novel causal inference framework used to quantitatively assess the external validity of both randomized and observational studies 1, 2, 3, 4, 5, 6. By evaluating external validity, transportability analyses help bridge the gap between effect estimates observed in a source population and their anticipated impact in the target population 7.
While different transportability methods are available, most require individual participant-level data (IPD) for both the source and target samples 3, 8, 9, 10, 11. Access to such data is often limited, presenting practical challenges to transportability analyses 1. A common scenario is that IPD is only available for the source, while the target population provides aggregate-level data (AgD), typically from disease registries, healthcare databases, or published studies. To address this limitation, recent methodological developments have sought to leverage AgD to improve the feasibility of transportability analyses. For instance, entropy balancing approaches use convex optimization to derive sampling weights that align covariate moments between IPD from the source sample and AgD from the target population, facilitating effect estimate transport. This approach has been refined with calibration techniques, enhancing robustness and accuracy across varied contexts, including randomized clinical trials (RCTs) and observational studies 12, 13. More recently, Quan et al. 14 introduced a double inverse probability weighting (DIPW) method tailored for situations with limited data from the target population, particularly when only regional summary data is available due to privacy constraints. Additionally, some methods incorporate more flexible weighting strategies that use individual-level covariates from the source sample, further improving estimation consistency and efficiency across broader scenarios 15.
While previous studies on these AgD-based transportability methods have explored binary and continuous outcomes, there has been limited exploration for survival analyses critical for oncology. Given the importance of accurately applying trial results to broader populations, recent work has explored different ways to transport survival outcomes from an RCT to target populations. For example, Ramagopalan et al. developed a method for transporting survival estimates: they transported overall survival (OS) effect estimates from US clinical practice to the Canadian population using baseline covariates adjustments based on the Canadian population 16. Zuo et al. applied the proportional hazards and accelerated failure time models to transport survival outcomes in adjuvant colon cancer trials 17. They explored one-trial-at-a-time and leave-one-trial-out methods, demonstrating how varying model structures and moment adjustments can address treatment effect discrepancies due to heterogeneous covariate distributions across trials, providing a structured framework for extending survival results to diverse populations. Lee et al. extended the augmented calibration weighting (ACW) methods with the linear spline-based hazard regression model to generalize treatment effects from the ACTG 175 HIV trial to various target populations without the limitation of proportional hazard (PH) assumption 18.
Notably, censoring in survival analysis critically impacts the performance of traditional methods and presents similar challenges for transportability. Recent efforts have focused on methods to address censoring effects specifically in survival data transportability. One such effort is Berkowitz et al.’s approach, which uses inverse odds of selection weights with survival tools like the Kaplan–Meier estimator and marginal structural Cox models to transport survival outcomes 19. Although effective for estimating counterfactual survival, this method assumes censoring is independent of covariates, limiting its real-world applicability where censoring often depends on covariates. Lee et al. advance this by addressing covariate-dependent censoring in survival transportability 20. Using a doubly robust (DR) estimator, their method integrates inverse probability weighting for sampling, censoring, and treatment assignment with outcome regression, ensuring accuracy even when censoring varies with covariates. Cao et al. expanded on weighting and DR methods to estimate counterfactual survival functions within the target population 21. Unlike Lee et al.’s calibration weighting estimators, this approach employs simpler estimators that rely on direct weight estimation or survival outcome modelling. However, it assumes that the target population data is available through a complex survey with pre-estimated weights, a more idealized scenario than purely aggregate-level data, as it requires well-estimated survey weights for accurate covariate alignment. While these methods have advanced survival outcome transportability, they still largely depend on complete or partial access to individual-level data from the target population, limiting their applicability when only aggregate-level data is available. This limitation highlights a gap in current research, motivating our efforts to enable reliable transportability with covariate-dependent censoring using limited target data.
In this paper, we propose a novel framework called “Target Aggregate Data Adjustment” (TADA) that can account for covariate-dependent censoring when transporting findings to aggregate-level target data. TADA adjusts for potential bias from censoring by incorporating the inverse probability of censoring weights and matches effect modifier distributions between source and target populations using the method of moments (MoM) weighting. TADA then derives final matching weights for each individual in the source dataset through a two-stage scheme, balancing the source and target populations to transport treatment effects.
The paper is organized as follows. In Section 2, we introduce the notations and formalize the causal inference framework for survival outcomes in transportability with necessary assumptions. Section 3 introduces the methodological approaches used in TADA. Sections 4 and 5 cover the simulation settings and present the results, respectively. A discussion is provided in Section 6, with concluding remarks in Section 7. A brief review of existing transportability methods for aggregate target data is provided in the Supplementary Materials.
2 Notation, Framework and Assumptions
Suppose we are interested in comparing the effectiveness of two treatments. We have access to the source sample $\mathcal{S}$ of size $n_{s}$, with individual data. Let $\textbf{X}\in\mathcal{X}$ be a vector of baseline covariates including effect modifiers for the individuals in the trial and those in the target population. $A\in\{0,1\}$ is the treatment indicator, with $A=1$ referring to active treatment and $A=0$ representing control. $S$ is defined as the binary trial participation indicator where $S=1$ stands for the individual participants in the RCT and $S=0$ otherwise. $T$ is the survival time, defined as the time of event occurrence. Following the potential outcomes framework 22, 23, let $T^{(a)}$ be the potential survival time if a subject receives the treatment $A=a$. Under the stable unit treatment value assumption (SUTVA) 22, the survival time is indicated as $T=T^{(1)}A+T^{(0)}(1-A)$. Let $C$ be the censoring time. In the presence of right censoring, the survival time $T_{i}$ is not always observed for all subjects; instead, we observe $U_{i}=T_{i}\wedge C_{i}$ and $\Delta=I(T\leq C)$, where $\wedge$ represents the minimum of two values, and $I(\cdot)$ is an indicator function. In summary, from the RCT source sample, we observe $\{U_{i},\Delta_{i},A_{i},\textbf{X}_{i},S_{i}=1\}$ from $i=1,\cdots,n_{s}$ subjects.
The target population, $\mathcal{T}$, contains $n_{t}$ individuals. We do not observe either the treatment or outcome but only have access to the sample moments $\bar{h}_{k,\mathcal{T}}=\frac{1}{n_{t}}\sum_{i\in\mathcal{T}}h_{k}(X_{i})$ of certain covariate functions $h_{k}:\mathcal{X}\rightarrow\mathbb{R}$, where $k=1,\dots,K_{h}$.
Let $S_{a}(t)=\Pr(T^{(a)}\geq t)$ be the survival function and $\lambda_{a}(t)=\text{lim}_{h\rightarrow 0}\frac{1}{h}\frac{\Pr(t\leq T^{(a)}\leq t+h)}{\Pr(T^{(a)}\geq t)}$ be the hazard function. Under the proportional hazards assumption, i.e., $\lambda_{1}(t)/\lambda_{0}(t)$ being a constant, the estimand of interest is chosen as the marginal hazard ratio, derived by estimating the causal effect of treatment on survival time within a weighted Cox proportional hazards (PH) model.
In RCTs, treatment assignment $A$ is independent of covariates X given study participation, thus the treatment propensity score $\pi_{A}$ is defined as the fixed probability of receiving the treatment. Let $\pi_{S}(\textbf{X})=\Pr(S=1\mid\textbf{X})$ be the sampling score. We make the following assumptions for the identification of the estimand of interest:
Assumption 1. (Randomization). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp A\mid S=1$; and $10<\pi_{A}<$ 1.
Assumption 2. (Ignorability and positivity of trial participation). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp S\mid\textbf{X}$; and $10<\pi_{S}(\textbf{X})<$ 1 for all values of X.
Assumption 3. (Noninformative censoring conditional on covariates and treatment). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$, which also implies $T\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$.
Given the inherent difficulties in directly validating these assumptions, relevant domain knowledge should support their practical plausibility. Assumption 1 holds in the RCT by design. If all information related to trial participation and outcomes is captured in the data, then the first part of Assumption 2 is reasonable. Additionally, the second part of Assumption 2 suggests that missing certain patient characteristics would prevent those patients from participating in the trial 24. Assumption 3 is the common assumption in survival analysis and imposes fewer restrictions compared to the conditional independence assumption between censoring and survival time based solely on treatment 8, 19, 25, 26.
3 Target Aggregate Data Adjustment (TADA)
3.1 Participation Weights: Method of Moments
If IPD is available for both the source population and target population, it is straightforward to obtain participation weights with regression-based methods. However, when IPD is not available for the target population, the limited information provided by AgD leads to additional challenges in estimation and requires a different approach. To address this issue, we propose a MoM-based approach as a feasible alternative for covariate adjustment.
The MoM method’s original application occurs in indirect treatment comparison (ITC), which also requires a similar covariate matching between two RCTs 27. As one of the earlier methods developed for ITC, MoM was initially applied in standard matching-adjusted indirect comparison (MAIC) and has since become a widely adopted approach in ITC. For each covariate of interest, the MoM approach aligns the sample moments calculated from the IPD of the source population with the corresponding moments from the AgD of the target population. Greater availability of AgD, particularly higher-order target sample moments, reduces bias in the population matching process. For instance, the access to additional summary statistics such as the sample standard deviation, median, proportion, or even higher-order moments (e.g., skewness or kurtosis) for covariates results in more accurate estimators compared to relying solely on the sample mean 28, 29. Our approach is conceptually similar to the strategy introduced by Josey et al. 12 where entropy balancing was employed to achieve the same goal. The existing literature highlights the connection between entropy balancing and the MoM approach, particularly in applications to ITC and transportability analyses. Both methods are statistically equivalent, sharing common advantages and robustness 30, while MoM has the advantage of being relatively intuitive in clinical interpretation, discussed in the Supplementary Materials.
Here, we illustrate the theoretical process of participation weights estimation by the MoM approach via first-order sample moments (i.e., sample mean) matching between two populations. Consider covariates X available in both source and target population. Assume that we have the availability of IPD for X in the source sample but just AgD for X in the target sample. For each individual $i$, let $X_{i}\in\textbf{X}_{i}$ be the available IPD in the source sample. Let $\bar{X}$ be the available AgD summary statistic in the target sample.
Assume that the weight for individual $i$ follows the logistic regression:
$w_{i}^{\textbf{MoM}}=\exp(\alpha+X_{i}\cdot k).$
Then $w_{i}^{\textbf{MoM}}$ can be estimated by matching $\bar{X}$ of the target sample with the weighted source sample moment as follows:
$\frac{\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}}{\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}}=\bar{X},$
where the coefficients $\alpha$ and $k$ in the weights will be estimated to satisfy the equality constraint.
Without loss of generality, it can be assumed that $\bar{X}=0$ in theoretical formula (e.g., we could transform/center baseline characteristics in both trials by subtracting $\bar{X}$ in format) and derive the procedure as follows:
$\frac{\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}}{\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}}$ $=\bar{X}$
$\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}$ $=\bar{X}\cdot\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}$
$\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}$ $=0=Q^{\prime}(k),$
where $Q(k)=\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}=\sum_{i=1}^{N}\exp(\alpha+X_{i}\cdot k)$. Since $Q(k)$ is convex, any solution of the equality constraint is unique and corresponds to the global minimum of $Q(k)$. We can follow the process to eventually get the estimates of $w_{i}^{\textbf{MoM}}$.
3.2 Censoring Weights: Inverse Probability of Censoring Weight
In longitudinal studies, participants are lost to follow-up for various reasons that may be related to the treatment or competing events. This loss to follow-up can result in systematic differences between those who remain in the study and those who do not, across treatment groups. While RCTs generally eliminate confounding through randomization, censoring can still introduce bias if the likelihood of being censored is related to the outcome the participant would have experienced in the absence of censoring. To address this issue, censoring weights are applied during the estimation process, aiming to recreate a hypothetical population that would have been observed without censoring. TADA aims to apply inverse probability censoring weights (IPCW) to account for bias resulting from censoring for the time-varying outcome in the source population.
Let $A=1$ represent the presence of treatment and $A=0$ represent the absence of treatment. Let $c$ be the censoring indicator where $c=1$ represents being censored and $c=0$ represents not being censored. Let X represent the baseline covariates related to censoring and the outcome. The standard inverse probability of censoring weight $w^{\textbf{c}}$ is the inverse of the conditional probability of not being censored given the baseline covariates and treatment status:
$w^{c}=\frac{1}{\Pr(c=0\mid\textbf{X},A)}.$
In an RCT, where patients are allocated to two parallel arms, the primary endpoint, such as overall survival, might be right-censored. In this case, at any given time $t$, the censoring weights are defined as the inverse of the probability that an individual remained on the originally randomized treatment (i.e., uncensored) until time $t$. The weights for each individual up to the censoring time can be defined as:
$w_{i}^{\textbf{c}}=\frac{1}{P(C_{i}>T_{i}\mid\textbf{X}=\textbf{x}_{i})},$
where $C_{i}$ is the time to censoring for individual $i$ and $T_{i}$ is the time to event for individual $i$.
To address the issue of unstable weights due to possibly extreme probability as denominators, stabilized weights are often used with the advantages of less variable and less skewed, as suggested by Robins 31. The stabilized censoring weights are defined as:
$sw_{i}^{\textbf{c}}=\frac{\bar{P}(C_{i}>T_{i})}{P(C_{i}>T_{i}\mid\textbf{X}=\textbf{x}_{i})},$
where the weight denominators remain constant, the numerators reflect the overall survival probability (the average probability of remaining uncensored).
We utilize the Weibull distribution to model both event times and censoring times, implying that the baseline hazard functions are time-varying 32, 33. The hazard function of the Weibull distribution is defined as:
$h(t)=\lambda\alpha t^{\alpha-1},$
where $\lambda$ is the scale parameter, $\alpha$ is the shape parameter, and $t$ represents time. When $\alpha\neq 1$, the hazard function varies with time: If $\alpha>1$, the hazard increases over time, If $\alpha<1$, the hazard decreases over time.
The cumulative hazard function is:
$H(t)=\lambda t^{\alpha},$
and the survival function is:
$S(t)=\exp\left\{-H(t)\right\}=\exp\left(-\lambda t^{\alpha}\right).$
We use the Cox PH model 34, 35 to model the relationship between censoring times and covariates:
$h_{c}(t\mid\textbf{X}_{i})=h_{c0}(t)\exp\left(\beta_{c}^{\top}\textbf{X}_{i}\right),$
where $h_{c}(t\mid X_{i})$ is the hazard function for censoring for individual $i$ at time $t$, $h_{c0}(t)$ is the time-varying baseline hazard function, $\beta_{c}$ is the vector of regression coefficients, and $\textbf{X}_{i}$ is the vector of covariates.
We estimate the baseline cumulative hazard function $H_{c0}(t)$ using the Breslow estimator 36:
[…]
where $d_{j}$ is the number of censoring events at time $t_{j}$, and $R(t_{j})$ is the risk set at time $t_{j}$.
For each individual $i$, we calculate the linear predictor:
$\eta_{i}=\beta_{c}^{\top}\textbf{X}_{i}.$
The probability that individual $i$ is uncensored at their observed time $t_{i}$ is:
$S_{c}(t_{i}\mid\textbf{X}_{i})=\exp\left\{-H_{c0}(t_{i})\exp\left(\eta_{i}\right)\right\}.$
The censoring weight $w_{i}$ for individual $i$ is defined as the inverse of their uncensored probability:
$w^{\textbf{c}}_{i}=\frac{1}{S_{c}(t_{i}\mid\textbf{X}_{i})},$
and the final stabilized censoring weights $sw_{i}^{\textbf{c}}$ are then used in TADA to adjust for bias due to censoring.
3.3 Final Weights: A Two-Staged Scheme
We introduce the final weights $w_{i}^{\textbf{final}}$ for each individual as the product of participation weights $w_{i}^{\textbf{MoM}}$ and the stabilized censoring weights $sw_{i}^{\textbf{c}}$ for all covariates of interest selected for population matching adjustment as follows:
$w_{i}^{\textbf{final}}=w_{i}^{\textbf{MoM}}\times sw_{i}^{\textbf{c}}.$
To mitigate the negative impact of extreme weights on estimation, we apply truncation at the 99% quantile. The final weights adjust for both the bias from censoring and the differences in effect modifier distributions between the source and target populations. This combined weighting approach ensures unbiased treatment effect estimates concerning both censoring and population differences. The weighted source population should closely resemble the target population, demonstrating population balance and supporting the treatment effect estimates.
In TADA, we employ a marginal structural model (MSM) to estimate the causal effect of treatment on survival outcomes. Since our weights are solely based on baseline covariates and do not involve any time-dependent components, we use a simplified MSM suitable for this context 31, 37. The MSM is specified as a weighted Cox PH model 34:
$\lambda_{a}(t\mid A_{i})=h_{0}(t)\exp(\beta A_{i}),$
where $h(t\mid A_{i})$ is the hazard function at time $t$ for individual $i$ with treatment assignment $A_{i}$, $h_{0}(t)$ is the baseline hazard function, and $\beta$ represents the log hazard ratio of the treatment effect.
We estimate $\beta$ by fitting the weighted Cox model using the final weights $w_{i}^{\textbf{final}}$ obtained from our two-stage weighting scheme as detailed in Sections 3.1 to 3.3. By applying these final weights in the Cox model we create a pseudo-population in which: the treatment is independent of baseline covariates; the bias due to censoring is reduced; and balance in effect modifiers between the source and target populations is achieved. This allows us to estimate the transported treatment effect on the target population using aggregate-level data.
The weighted partial likelihood function for the Cox model 38 is given by:
$L(\beta)=\prod_{i=1}^{n_{s}}\left[\frac{\exp(\beta A_{i})}{\sum_{j\in R(t_{i})}w_{j}^{\textbf{final}}\exp(\beta A_{j})}\right]^{\delta_{i}w_{i}^{\textbf{final}}},$
where $n_{s}$ is the number of individuals in the source sample, $\delta_{i}$ is the event indicator for individual $i$, and $R(t_{i})$ is the risk set at time $t_{i}$. Maximizing this weighted likelihood function provides an estimate of $\beta$, the log hazard ratio, which quantifies the causal effect of the treatment in the target population. We are estimating the variance of $\beta$ using the bootstrap method.
By integrating the MSM into TADA, the treatment effect estimator is unbiased concerning both censoring and population imbalance. The weighted source sample is adjusted to closely resemble the target population in terms of aggregate-level summaries, thereby enhancing the transportability of causal inferences.
4 Simulation Study
The simulation study applies the ADEMP framework proposed by Morris et al. 39 to comprehensively evaluate the effectiveness of TADA in transportability analysis using AgD for a target population, specifically examining its performance in adjusting for censoring. Our primary goal (Aim) is to assess the estimated causal effects of treatment on overall survival across different RCT scenarios, with and without censoring adjustments made via TADA. Detailed descriptions of the Data-generating mechanisms, Estimands, Methods, and Performance measures are provided in the following subsections.
4.1 Data-Generating Mechanism
In each simulation replicate, we generate three datasets named studyData, targetData, and targetData_Popula-tion_IPD respectively: studyData contains simulated individual-level data from the source population, including treatment status, covariates, event times, censoring times, observed times, and indicators for observed and censored events. targetData provides aggregate summary statistics (proportions, means, and standard deviations of each covariate) for the target population, matching those in studyData. targetData_Population_IPD consists of pseudo individual-level data for the target population, including covariates, treatment status, and event times.
To generate studyData, we first simulate a large source population of 50,000 individuals and then randomly sampled 200 individuals for the final dataset. Each individual has three baseline covariates: $X_{1}$ and $X_{2}$ are binary variables generated from binomial distributions with parameters 0.45 and 0.65, respectively. $X_{3}$ is a continuous variable generated from a standard normal distribution (mean 0, variance 1). Treatment status is assigned randomly to simulate an RCT scenario, with an allocation rate of 1:1.
Event times are generated from a Weibull distribution, incorporating the effects of covariates and treatment through the linear predictor $\eta_{\text{event},i}$.
$\eta_{\text{event},i}$ $=\beta_{X_{1}}X_{1,i}+\beta_{X_{2}}X_{2,i}+\beta_{X_{3}}X_{3,i}+\beta_{\text{trt}}\cdot\text{Treatment}_{i}$
$+\beta_{\text{trt-$X_{1}$}}\cdot\text{Treatment}_{i}\cdot X_{1,i}+\beta_{\text{trt-$X_{2}$}}\cdot\text{Treatment}_{i}\cdot X_{2,i}+\beta_{\text{trt-$X_{3}$}}\cdot\text{Treatment}_{i}\cdot X_{3,i},$
where the main effects are set as $\beta_{X_{1}}=0.5$, $\beta_{X_{2}}=-0.3$, $\beta_{X_{3}}=0.2$, and $\beta_{\text{trt}}=-0.4$.
We generate a total of 10 distinct random scenarios by varying the coefficient combinations of the treatment-covariate interaction terms $(\beta_{\text{trt-$X_{1}$}},\beta_{\text{trt-$X_{2}$}},\beta_{\text{trt-$X_{3}$}})$, while keeping the main effect coefficients $\beta_{X_{1}}$, $\beta_{X_{2}}$, $\beta_{X_{3}}$ and $\beta_{\text{trt}}$ constant. This approach ensures that all variables are sufficiently perturbed during data generation, creating a range of representative and complex simulation scenarios. The coefficient combinations are provided in Table 1 by scenarios.
Scenario $\beta_{\text{trt-$X_{1}$}}$ $\beta_{\text{trt-$X_{2}$}}$ $\beta_{\text{trt-$X_{3}$}}$
1 1.20 0.35 1.30
2 0.58 0.35 1.30
3 -0.90 0.35 1.30
4 -0.60 0.35 1.30
5 1.20 0.20 1.30
6 1.20 -0.70 1.30
7 1.20 -0.10 1.30
8 1.20 0.35 0.45
9 1.20 0.35 -0.30
10 1.20 0.35 -0.85
Table 1: 10 distinct random scenarios and corresponding coefficient combinations of the treatment-covariate interaction terms $(\beta_{\text{trt-$X_{1}$}},\beta_{\text{trt-$X_{2}$}},\beta_{\text{trt-$X_{3}$}})$.
Event times $T_{i}$ are simulated to model time-to-event data with a time-varying baseline hazard function using the inverse cumulative distribution function (CDF) of Weibull distribution:
$T_{i}=\left(-\frac{\ln(u_{i})}{\lambda_{\text{event}}\exp\left(\eta_{\text{event},i}\right)}\right)^{1/\alpha_{\text{event}}},$
where shape parameter $\alpha_{\text{event}}=1.5$ and scale parameter $\lambda_{\text{event}}=0.1$. These parameters are chosen to reflect an increasing hazard over time ($\alpha_{\text{event}}>1$). $u_{i}$ is a random variable drawn from the uniform distribution on (0, 1).
In studyData, we generate censoring time for each individual following a Weibull distribution with the impact of both treatment and covariates. A linear predictor $\eta_{\text{censor},i}$ is constructed for each individual to incorporate the effects of covariates on the hazard function.
$\eta_{\text{censor},i}=\beta_{0}+\beta_{X_{1}}^{c}X_{1,i}+\beta_{X_{2}}^{c}X_{2,i}+\beta_{X_{3}}^{c}X_{3,i}+\beta_{\text{trt}}^{c}\cdot\text{Treatment}_{i},$
where $\beta_{X_{1}}^{c}=-0.2$, $\beta_{X_{2}}^{c}=0.4$ and $\beta_{X_{3}}^{c}=-0.1$ as the main effect of baseline covariates, $\beta_{\text{trt}}^{c}=0.25$ as the treatment main effect and $\beta_{0}$ as the intercept. We choose various candidates of $\beta_{0}$ to directly adjust the overall censoring proportion among studyData. We set $\beta_{0}$ as 2.5, 3.3, 3.7 and 4.3 respectively to obtain studyData with around 20%, 30%, 40% and 50% overall censoring proportion while keeping other coefficients being consistent.
The censoring times $C_{i}$ are generated using:
$C_{i}=\left(-\frac{\ln(u_{i}^{\prime})}{\lambda_{\text{censor}}\exp\left(\eta_{\text{censor},i}\right)}\right)^{1/\alpha_{\text{censor}}},$
where shape parameter $\alpha_{\text{censor}}=1.5$ and scale parameter $\lambda_{\text{censor}}=0.001$. $u_{i}^{\prime}$ is an independent random variable from the uniform distribution on (0, 1).
For each individual, the observed time ($U_{i}$) is determined by:
$U_{i}=T_{i}\wedge C_{i},$
where $\wedge$ represents the minimum of two values.
The event indicator $\Delta_{i}$ is defined as:
$\Delta_{i}=\begin{cases}1,&\text{if }T_{i}\leq C_{i}\quad(\text{event observed}),\\ 0,&\text{if }T_{i}>C_{i}\quad(\text{censored}).\end{cases}$
Similarly, we generate targetData by simulating a large target population of 200,000 individuals and then randomly sampling 1,000 individuals for the final dataset. Each individual has three baseline covariates as same as in the source population ($X_{1}$, $X_{2}$, and $X_{3}$) with different distribution characteristics: $X_{1}$ is binary and generated from a binomial distribution with parameter 0.35. $X_{2}$ is binary and generated from a binomial distribution with parameter 0.55. $X_{3}$ is continuous and generated from a normal distribution with a mean of 0.21 and a variance of 1.5. Since targetData is an aggregate-level dataset, we summarize the proportions of $X_{1}$ and $X_{2}$, and the mean and standard deviation of $X_{3}$ as the final data.
The individual-level covariate data for targetData_Population_IPD are generated concurrently before aggregation. To obtain the “true” treatment effect in the target population for benchmarking, we assign treatment status and generate event times for each individual in targetData_Population_IPD. Treatment status is randomly assigned with a 50% probability to simulate an RCT scenario. We use the same Weibull distribution parameters ($\alpha_{\text{event}}$, $\lambda_{\text{event}}$), the same linear predictor structure $\eta_{\text{event},i}$ and coefficients as in studyData, applying them to targetData_Population_IPD to ensure consistency.
4.2 Estimands, Targets and Methods
We are interested in estimating the marginal hazard ratio in the target population. We consider two analytical strategies for handling censoring in the context of transportability analysis.
  • Adjusting censoring: In this case, TADA adjusts the bias resulting from the presence of censoring via obtaining the IPCW introduced in Section 3.1, as well as balancing the source and target population with the participation weights introduced in Section 3.2. The final weights are as introduced in Section 3.3 comprised of both weights.
  • Ignoring censoring: In this case, TADA ignores the impact of the presence of censoring and only balances the source and target population with the participation weights introduced in Section 3.2. The final weights are equivalent to the participation weights.
For both strategies, we estimate average marginal hazard ratios using a Cox PH model based on multiple replicates. Meanwhile, we apply the Cox PH model on targetData_Population_IPD to estimate average marginal hazard ratios as the pseudo “true HR” of the target population as the comparison benchmark.
4.3 Performance Measures
The performance measures of interest are the bias of the estimators calculated using the above methods and their coverage of the 95% confidence intervals compared to the analysis in the IPD target data. Let $\hat{\gamma}_{\text{cen,i}}$ be the estimated marginal hazard ratios for the $i=1,2,\cdots,N$-th simulated replicate when considering the adjustment of censoring. Let $\hat{\gamma}_{\text{non-cen,i}}$ be the estimated marginal hazard ratios for the $i=1,2,\cdots,N$-th simulated replicate when ignoring the adjustment of censoring. With the pseudo “true HR” of target population being $\bar{\gamma}$, the bias of two cases aforementioned are estimated with
$\widehat{\text{Bias}}_{\text{cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{\text{cen,i}}-\bar{\gamma},$
$\widehat{\text{Bias}}_{\text{non-cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{\text{non-cen,i}}-\bar{\gamma},$
and the coverages are estimated with
$\widehat{\text{Coverage}}_{\text{cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}I\left\{\hat{\gamma}_{\text{cen,i}}-1.96\hat{se}(\gamma)_{\text{cen,i}}\leq\bar{\gamma}\leq\hat{\gamma}_{\text{cen,i}}+1.96\hat{se}(\gamma)_{\text{cen,i}}\right\},$
$\widehat{\text{Coverage}}_{\text{non-cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}I\left\{\hat{\gamma}_{\text{non-cen,i}}-1.96\hat{se}(\gamma)_{\text{non-cen,i}}\leq\bar{\gamma}\leq\hat{\gamma}_{\text{non-cen,i}}+1.96\hat{se}(\gamma)_{\text{non-cen,i}}\right\}.$
We adopt the bootstrapping method to obtain the $\hat{se}(\gamma)_{\text{cen,i}}$ and $\hat{se}(\gamma)_{\text{non-cen,i}}$ with $B=1000$ sampling for each replicate. We simulate 1000 replicates for each case.
5 Results
In simulation study, we have evaluated the performance of TADA under different strategies for censoring. Tables 3 to 5 summarize key simulation results for both the censoring-adjusted and censoring-ignored approaches across all scenarios under various overall censoring proportions. Specifically, each table compares the estimated bias and coverage relative to the pseudo true hazard ratio. Figure 1 provides the distribution of the overall censoring rate of each scenario under 20% and 30% average overall censoring, and the censoring rates of 40% and 50% for Figure 2. Figure 3 presents the censoring rates for the treatment and control groups across various scenarios under overall censoring rates of 20% and 30%, while Figure 4 illustrates the censoring rates in each group across scenarios under average overall censoring rates of 40% and 50%.
In general, TADA with censoring adjustments demonstrates superior performance in terms of bias and coverage compared to the unadjusted approach across various censoring scenarios. As the difference in censoring rates between treatment and control groups grows (e.g., in Scenarios 3, 4, 5 and 6), the benefits of the censoring-adjusted method become more evident across overall censoring rates ranging from 20% to 50%. Compared to the true hazard ratio, the censoring-adjusted TADA consistently yields lower bias and improved coverage in most cases.
At lower overall censoring rates (20% and 30%), the advantages of the censoring-adjusted TADA are particularly prominent. For instance, under a 20% censoring rate, the censoring-adjusted hazard ratio estimates show a marked reduction in bias compared to the unadjusted estimates—such as in Scenario 1, where the adjusted bias is 0.037 versus 0.128 for the unadjusted approach, roughly one-fourth of the unadjusted bias. Significant improvements are seen in Scenarios 3 and 4, with censoring-adjusted bias estimates close to zero compared to more than twenty times the results without adjustment. Similarly, in Scenario 6, the censoring-adjusted bias decreases from 0.085 to 0.003. Additionally, the censoring-adjusted method achieves coverage exceeding 92% in most scenarios at these lower censoring rates, indicating robust performance.
As the overall censoring rate rises to 40%, the censoring-adjusted TADA continues to demonstrate lower bias than the unadjusted approach, though both methods show some decline in performance. In Scenario 3, for example, the adjusted bias is 0.139 compared to an unadjusted bias of 0.218, both exceeding 10%. However, the coverage of the censoring-adjusted estimates remains higher than that of the unadjusted estimates, achieving 81.6% versus 77.1% in Scenario 3, which underscores the robustness of the adjusted method under moderate censoring conditions.
At an overall censoring rate of 50%, the benefits of censoring adjustment, though somewhat reduced, are still present. For instance, in Scenario 9, the censoring-adjusted bias is 0.190 compared to 0.196 for the unadjusted approach. While both methods experience coverage reductions at this level, the censoring-adjusted TADA maintains a higher coverage rate, such as 80.3% versus 78.5% in Scenario 8, demonstrating its relative resilience even under more challenging censoring conditions.
As shown in Figure 3 and 4, in all scenarios, the censoring proportions between the treatment and control groups do not exhibit extreme discrepancies, meaning there is no scenario where one group had a very low censoring rate while the other had a very high rate. This balanced design across scenarios ensures that the adjustment results from TADA are broadly applicable and robust.
Figure 1: The boxplots of overall censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from left to right represent the minimum, mean and maximum of the overall censoring proportion among 1000 simulation replicates.
Figure 2: The boxplots of overall censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 40% and (B) 50%. The numbers aligned with each boxplot from left to right represent the minimum, mean and maximum of the overall censoring proportion among 1000 simulation replicates.
Figure 3: The boxplots of treatment (bright orange) and control (light blue) group censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from bottom to top represent the minimum (blue), mean (black) and maximum (red) of corresponding censoring proportion among 1000 simulation replicates.
Figure 4: The boxplots of treatment (bright orange) and control (light blue) group censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 40% and (B) 50%. The numbers aligned with each boxplot from bottom to top represent the minimum (blue), mean (black) and maximum (red) of corresponding censoring proportion among 1000 simulation replicates.
Scenario HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
1 0.855 (0.619, 1.435) 0.946 (0.627, 1.440) 0.818 (0.809, 0.828) 0.037 (-0.266, 0.433) 0.128 (-0.188, 0.526) 0.926 0.925
2 0.774 (0.559, 1.304) 0.862 (0.569, 1.312) 0.753 (0.743, 0.762) 0.021 (-0.251, 0.387) 0.109 (-0.189, 0.506) 0.925 0.918
3 0.587 (0.425, 1.009) 0.666 (0.436, 1.020) 0.591 (0.582, 0.601) -0.004 (-0.220, 0.263) 0.075 (-0.150, 0.367) 0.935 0.923
4 0.620 (0.448, 1.062) 0.700 (0.458, 1.072) 0.623 (0.613, 0.633) -0.003 (-0.247, 0.296) 0.077 (-0.184, 0.391) 0.943 0.932
5 0.815 (0.592, 1.374) 0.907 (0.600, 1.380) 0.788 (0.779, 0.798) 0.026 (-0.273, 0.405) 0.119 (-0.189, 0.505) 0.924 0.916
6 0.626 (0.454, 1.070) 0.709 (0.465, 1.082) 0.624 (0.614, 0.633) 0.003 (-0.234, 0.318) 0.085 (-0.167, 0.420) 0.929 0.922
7 0.749 (0.545, 1.271) 0.840 (0.555, 1.277) 0.730 (0.721, 0.740) 0.019 (-0.243, 0.381) 0.110 (-0.167, 0.489) 0.924 0.917
8 0.989 (0.706, 1.587) 1.048 (0.709, 1.585) 0.952 (0.943, 0.961) 0.037 (-0.302, 0.461) 0.097 (-0.247, 0.522) 0.938 0.940
9 1.032 (0.725, 1.589) 1.065 (0.725, 1.588) 0.990 (0.982, 0.999) 0.042 (-0.306, 0.506) 0.075 (-0.267, 0.515) 0.950 0.947
10 0.901 (0.636, 1.415) 0.946 (0.639, 1.415) 0.796 (0.787, 0.805) 0.105 (-0.196, 0.527) 0.149 (-0.153, 0.554) 0.873 0.869
Table 2: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 20%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
Scenario HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
1 0.973 (0.683, 1.687) 1.075 (0.698, 1.693) 0.818 (0.809, 0.828) 0.157 (-0.208, 0.656) 0.257 (-0.113, 0.714) 0.814 0.783
2 0.882 (0.613, 1.527) 0.975 (0.629, 1.538) 0.753 (0.743, 0.762) 0.129 (-0.215, 0.593) 0.222 (-0.137, 0.681) 0.833 0.800
3 0.669 (0.459, 1.180) 0.748 (0.478, 1.195) 0.591 (0.582, 0.601) 0.078 (-0.188, 0.450) 0.157 (-0.126, 0.524) 0.865 0.841
4 0.704 (0.487, 1.241) 0.788 (0.505, 1.257) 0.623 (0.613, 0.633) 0.081 (-0.192, 0.441) 0.165 (-0.138, 0.538) 0.860 0.833
5 0.931 (0.651, 1.612) 1.028 (0.667, 1.619) 0.788 (0.779, 0.798) 0.142 (-0.208, 0.641) 0.240 (-0.123, 0.719) 0.824 0.791
6 0.726 (0.500, 1.268) 0.810 (0.519, 1.284) 0.624 (0.614, 0.633) 0.102 (-0.199, 0.487) 0.186 (-0.126, 0.584) 0.834 0.804
7 0.863 (0.602, 1.499) 0.957 (0.620, 1.508) 0.730 (0.721, 0.740) 0.133 (-0.184, 0.595) 0.227 (-0.101, 0.655) 0.819 0.795
8 1.072 (0.733, 1.760) 1.125 (0.741, 1.755) 0.952 (0.943, 0.961) 0.120 (-0.286, 0.630) 0.173 (-0.226, 0.694) 0.906 0.897
9 1.059 (0.719, 1.695) 1.095 (0.724, 1.689) 0.990 (0.982, 0.999) 0.068 (-0.306, 0.616) 0.104 (-0.269, 0.612) 0.940 0.936
10 0.951 (0.647, 1.540) 0.996 (0.656, 1.536) 0.796 (0.787, 0.805) 0.155 (-0.196, 0.657) 0.200 (-0.164, 0.693) 0.845 0.837
Table 3: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 30%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
Scenario HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
1 1.067 (0.731, 1.890) 1.174 (0.751, 1.896) 0.818 (0.809, 0.828) 0.249 (-0.176, 0.830) 0.356 (-0.067, 0.914) 0.708 0.680
2 0.962 (0.652, 1.705) 1.059 (0.673, 1.713) 0.753 (0.743, 0.762) 0.209 (-0.174, 0.728) 0.306 (-0.082, 0.844) 0.749 0.711
3 0.730 (0.487, 1.314) 0.810 (0.508, 1.331) 0.591 (0.582, 0.601) 0.139 (-0.150, 0.558) 0.218 (-0.086, 0.666) 0.816 0.771
4 0.770 (0.516, 1.385) 0.854 (0.537, 1.399) 0.623 (0.613, 0.633) 0.147 (-0.171, 0.563) 0.231 (-0.108, 0.649) 0.791 0.744
5 1.024 (0.697, 1.806) 1.124 (0.718, 1.813) 0.788 (0.779, 0.798) 0.236 (-0.174, 0.807) 0.336 (-0.083, 0.892) 0.724 0.690
6 0.804 (0.535, 1.428) 0.886 (0.557, 1.442) 0.624 (0.614, 0.633) 0.181 (-0.166, 0.642) 0.262 (-0.085, 0.714) 0.763 0.709
7 0.952 (0.644, 1.680) 1.045 (0.665, 1.690) 0.730 (0.721, 0.740) 0.222 (-0.167, 0.757) 0.315 (-0.069, 0.844) 0.737 0.692
8 1.136 (0.750, 1.895) 1.180 (0.763, 1.886) 0.952 (0.943, 0.961) 0.184 (-0.257, 0.842) 0.229 (-0.202, 0.812) 0.868 0.862
9 1.098 (0.718, 1.791) 1.124 (0.726, 1.780) 0.990 (0.982, 0.999) 0.108 (-0.316, 0.753) 0.134 (-0.270, 0.712) 0.928 0.923
10 0.993 (0.654, 1.640) 1.033 (0.665, 1.635) 0.796 (0.787, 0.805) 0.197 (-0.199, 0.767) 0.237 (-0.159, 0.776) 0.816 0.805
Table 4: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 40%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
Scenario HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
1 1.269 (0.832, 2.358) 1.386 (0.860, 2.353) 0.818 (0.809, 0.828) 0.451 (-0.093, 1.239) 0.567 (0.020, 1.325) 0.520 0.473
2 1.135 (0.738, 2.114) 1.240 (0.763, 2.112) 0.753 (0.743, 0.762) 0.382 (-0.099, 1.115) 0.487 (0.010, 1.191) 0.576 0.526
3 0.868 (0.543, 1.636) 0.945 (0.570, 1.647) 0.591 (0.582, 0.601) 0.277 (-0.117, 0.831) 0.354 (-0.034, 0.909) 0.668 0.606
4 0.911 (0.574, 1.711) 0.991 (0.600, 1.716) 0.623 (0.613, 0.633) 0.288 (-0.132, 0.932) 0.368 (-0.053, 0.961) 0.665 0.598
5 1.220 (0.795, 2.259) 1.330 (0.822, 2.253) 0.788 (0.779, 0.798) 0.432 (-0.082, 1.216) 0.542 (0.013, 1.293) 0.521 0.476
6 0.964 (0.605, 1.787) 1.044 (0.631, 1.791) 0.624 (0.614, 0.633) 0.340 (-0.106, 0.982) 0.420 (-0.000, 1.043) 0.582 0.523
7 1.135 (0.732, 2.108) 1.235 (0.760, 2.102) 0.730 (0.721, 0.740) 0.404 (-0.099, 1.143) 0.505 (0.017, 1.191) 0.530 0.481
8 1.283 (0.787, 2.221) 1.301 (0.804, 2.196) 0.952 (0.943, 0.961) 0.331 (-0.209, 1.207) 0.349 (-0.148, 1.087) 0.803 0.785
9 1.180 (0.714, 2.014) 1.186 (0.727, 1.992) 0.990 (0.982, 0.999) 0.190 (-0.301, 0.959) 0.196 (-0.243, 0.859) 0.912 0.902
10 1.083 (0.666, 1.878) 1.113 (0.683, 1.859) 0.796 (0.787, 0.805) 0.287 (-0.193, 0.995) 0.317 (-0.138, 0.931) 0.784 0.757
Table 5: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 50%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
6 Discussion
In this paper, we demonstrated that the proposed TADA method provides substantial improvements in bias control and coverage across a range of censoring scenarios, showing particular effectiveness under low to moderate censoring proportions. These findings support the value of TADA in addressing censoring bias, with its advantage over unadjusted methods persisting even as overall censoring increases, albeit with some reduction in performance. Overall, these results suggest that TADA structurally addresses information loss due to censoring, making it a robust transportability analysis method in scenarios where IPDs are unavailable for the target population and where source data are disturbed by right-censoring.
A core strength of the proposed TADA method lies in its flexible weighting by a product structure, with each term specifically designed to adjust for different sources of confounding. This makes TADA highly adaptable to complex application scenarios. In current cases in TADA, the participation weights, derived via the MoM, can address the challenges of working with AgD for the target population for transportability analysis. In this study, we explored TADA’s ability to transport survival outcomes under covariant-relevant censoring, which addresses the current gap in the transportability methods that generally have been limited to binary and continuous outcomes with requirements of IPD.
In scenarios with modest censoring rates, the TADA method with an adjustment for censoring exhibits robust performance with low bias and high coverage. This effectiveness can be attributed to the lower level of information loss in low-censoring contexts, allowing TADA’s inverse probability weighting to make more precise adjustments. The stability of TADA’s performance in these scenarios highlights its ability to achieve accurate effect estimates under conditions of limited censoring, where the method leverages available data efficiently. In scenarios with a higher overall censoring proportion, TADA demonstrates performance close to its practical limit, with a noticeable decline in bias control and coverage. Increased censoring results in a higher loss of information that in turn effectively reduces the sample size available for reliable adjustments. Higher censoring may result in more extreme weights and introduce greater variability in bias control. This accumulation of censoring-induced bias and weight instability creates challenges for maintaining accuracy and coverage under high-censoring conditions. Nonetheless, censoring-adjusted TADA outperforms the unadjusted approach in general.
In scenarios where there was a larger disparity in censoring proportions between treatment and control groups, TADA shows lower bias estimates. This phenomenon suggests that TADA may potentially benefit from this unique data characteristic. When one group has lower censoring, the more complete data could allow TADA’s weighting mechanism to draw on relatively stable information, which may enhance the reliability of adjustments. This natural heterogeneity might provide a form of implicit stratification, helping TADA balance distinct group effects with greater robustness. Such conditions could, in theory, offer additional stability in TADA’s performance by leveraging the more consistent information from less-censored data. However, as overall censoring increases, this differential effect diminishes, indicating that higher censoring rates may challenge TADA’s capacity to control bias across varying group censoring proportions.
Future research could expand the scope of TADA’s applicability by exploring alternative approaches to deriving weights for population balance and handling censoring adjustments. In the current framework, the MoM is employed to achieve population balance, while IPCW is used to adjust for censoring. However, the effectiveness of IPCW is somewhat reliant on the PH assumption, which may limit its utility in settings where this assumption does not hold. Exploring alternative censoring adjustment techniques that do not depend on the PH assumption could enhance TADA’s versatility in handling varied censoring structures 18. By experimenting with different weighting schemes for censoring adjustments, TADA could potentially expand its applicability in contexts where data characteristics or censoring patterns diverge from the ideal settings. This evolution would allow TADA to maintain its effectiveness under diverse data conditions, making it a more adaptable tool across a broad range of practical applications. Additionally, while TADA is currently applied for RCT source data, extending its framework to accommodate observational data could further broaden its utility, enabling transportability analyses in settings where randomized data is unavailable.
7 Conclusion
Our study demonstrates the TADA’s robust ability to effectively address censoring bias across a broad spectrum of scenarios, demonstrating superior bias control and coverage even under challenging conditions. The TADA method’s design for using aggregate data for causal inference, along with its ability to accommodate varying censoring proportions, makes it an invaluable tool for transporting survival estimates, especially in settings where target IPD is hard to access.
Acknowledgments
Conflict of Interest: None declared.